# Exponential Dichotomies in Higher-Dimensional Spatial Dynamics for Elliptic Partial Differential Equations

### v1

@@ 1 Introduction
We say that a radial spatial evolutionary equation of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy in a Banach space $Y$ on a sub-interval $J\subseteq\mathbb{R}^{+}$ if there exist positive constants $C$ and $\eta$, two strongly continuous family of projections $P^{\mathrm{s}/\mathrm{u}}:J\rightarrow L(Y)$ with $P^{\mathrm{s}}(r)+P^{\mathrm{u}}(r)=I$, and two families of linear operators $\Phi^{\mathrm{s}/\mathrm{u}}(r,\rho)$ such that the following statements are true:
  • (Stability) For all $\rho\in J$ and $\mathbf{u}_{0}\in Y$ there exists a solution $\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\geq\rho$, and differentiable in both $r$ and $\rho$ for $r>\rho$. The operator $P^{\mathrm{s}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{s}}(\rho,\rho)\mathbf{u}_{0}$ gives the stable projection, and
    $\|\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\|_{Y}\leq Ce^{-\eta(r-\rho)}\|\mathbf{u}_{0}\|_{Y},\qquad r\geq\rho,\,\,r,\rho\in J.$ (10)
  • (Instability) For all $\rho\in J$ and $\mathbf{u}_{0}\in Y$ there exists a solution $\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}$, defined and continuous for both $r$ and $\rho$ on $r\leq\rho$, and differentiable in both $r$ and $\rho$ for $r<\rho$. The operator $P^{\mathrm{u}}(\rho)\mathbf{u}_{0}:=\Phi^{\mathrm{u}}(\rho,\rho)\mathbf{u}_{0}$ gives the unstable projection, and
    $\|\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\|_{Y}\leq Ce^{\eta(r-\rho)}\|\mathbf{u}_{0}\|_{Y},\qquad r\leq\rho,\,\,r,\rho\in J.$ (11)
@@
    $~P^{\mathrm{s}}(r)\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)P^{\mathrm{s}}(\rho)$
@@
Since $X_{r}$ is a family of Banach spaces whose norm changes with $r$, we now define an analogous understanding of an exponential dichotomy on $X_{r}$.
Definition 2.2.
(Definition 5.2 in [17]) We say that a radial spatial evolutionary equation of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy in the $r$-dependent family of Banach spaces $X_{r}$ on a sub-interval $J\subseteq\mathbb{R}^{+}$ if there exist positive constants $C$ and $\eta$, two strongly continuous family of projections $P^{\mathrm{s}/\mathrm{u}}(r)\in L(X_{r})$ with $P^{\mathrm{s}}(r)+P^{\mathrm{u}}(r)=I$, and two families of linear operators $\Phi^{\mathrm{s}/\mathrm{u}}(r,\rho)\in L(X_{\rho},X_{r})$ such that the following statements are true:
    $\|\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq Ce^{-\eta(r-\rho)}\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\geq\rho,\,\,r,\rho\in J.$ (12)
    $\|\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq Ce^{\eta(r-\rho)}\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\leq\rho,\,\,r,\rho\in J.$ (13)
    $~P^{\mathrm{s}}(r)\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)P^{\mathrm{s}}(\rho)$
For completeness, we now also state the standard robustness result for exponential dichotomies. This result was originally stated for an arbitrary fixed infinite dimensional function space $Y$ in [11], however, the statement holds almost identically for the family of spaces $X_{r}$ [17]. Here we summarize both robustness statements in one.
@@
(Theorem 1 in [11] and Proposition 5.9 in [17]) Exponential dichotomies are robust under sufficiently small perturbations. More precisely, if a system of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy on some interval $J\subseteq\mathbb{R}^{+}$ in $Y$ (or in the $r$-dependent family of $X_{r}$ spaces) with constants $C,\eta>0$, then for all $\varepsilon>0$ and $\tilde{\eta}\in(0,\eta)$ there are constants $\tilde{C},\delta>0$ such that the perturbed system
@@
$\|\mathcal{B}(r)\|_{L(Y)}\leq\delta,\quad\qquad\big(\text{ or }\quad\|\mathcal{B}(r)\|_{L(X_{r})}\leq\delta,\big)\qquad r\in J$
@@
has an exponential dichotomy on $J$ in the same (family of) function space(s) with constants $\tilde{C}$ and $\tilde{\eta}$ and with projections $\varepsilon$-close to the projections of $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$. Furthermore, if $\|\mathcal{B}(r)\|_{L(Y)}=O(1/r)$ (or $\|\mathcal{B}(r)\|_{L(X_{r})}=O(1/r)$), then one can take $\tilde{\eta}=\eta$.
With these robustness results in hand, our aim is to prove the existence of an exponential dichotomy for the principal system (9) given the existence of an exponential dichotomy for the asymptotic system (6). Assuming that $V(r,\varphi)$ decays to $V_{\infty}$ sufficiently fast, Proposition 2.3 can then be applied directly to establish our main result. We therefore make the following hypothesis:
(H1.) $V(r,\varphi)$ decays to $V_{\infty}$ sufficiently fast as $r\to\infty$. That is, $\|V(r,\varphi)-V_{\infty}\|\leq C/r$ for large $r$.
@@
Since $\mathcal{A}_{\infty}$ is autonomous and homogeneous across $S^{n-1}$, the asymptotic system (6) has an exponential dichotomy whenever $\mathcal{A}_{\infty}$ is hyperbolic as a $2d\times 2d$ matrix. We can construct an exponential dichotomy for (6) via a finite spectral decomposition whenever the eigenvalues of $\mathcal{A}_{\infty}$ are bounded away from $i\mathbb{R}\subset\mathbb{C}$ or, equivalently, whenever the eigenvalues of $V_{\infty}$ are bounded away from the negative real line, $(-\infty,0]$. To use this condition later, we now set it aside in a hypothesis:
(H2.) All eigenvalues of $V_{\infty}$ are bounded away from the negative real line $(-\infty,0]$.
@@
Finally, we also assume that $V_{\infty}$ has only semi-simple eigenvalues, and is therefore diagonalizable. This last hypothesis greatly simplifies the proofs to follow, allowing for an explicit spectral decomposition of $\tilde{\mathcal{A}}(r)$ into scalar pieces as we will see in §3.1. However, similar arguments should hold in the case of nontrivial Jordan blocks.
@@
(H3.) $V_{\infty}$ is diagonalizable with eigenvalues $\{\lambda_{\ell}\}_{\ell=1}^{d}$ and corresponding eigenvectors $\{\rho_{\ell}\}_{\ell=1}^{d}\subset\mathbb{C}^{d}$.
@@
Hypotheses (H2) and (H3) allow us to fully characterize an exponential dichotomy for the asymptotic system (6). We must then understand the relationship between the principal operator and the asymptotic operator. To do this, we establish an isomorphism, or rather a family of isomorphisms, that can translate between the $X_{r}$ setting and the $L^{2}\!\!\times\!L^{2}$ setting. As we will see in Lemma 2.4, the basis of this family of isomorphisms is the differential operator defined by
@@
    The family of operators $A(r)^{1/2}:H^{1}(S^{n-1})\rightarrow L^{2}(S^{n-1})$ is well-defined, bounded uniformly in $r$, strongly continuous, and continuously differentiable in $r$, with
@@
    $C_{1}\left(\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}\right)\leq\|\mathbf{u}\|_{X_{r}}^{2}\leq C_{2}\left(\|A(r)^{1/2}u\|^{2}_{L^{2}}+\|v\|_{L^{2}}^{2}\right)$ (15)
@@
Lemma 2.4 establishes $(u,v)\mapsto(A(r)^{1/2}u,v)$ as an isomorphism from $X_{r}$ to $L^{2}\!\!\times\!L^{2}$. Applying the corresponding change of variables $\tilde{u}=A(r)^{1/2}u$ to the principal equation (9) we get the following equation posed on $L^{2}\!\!\times\!L^{2}$:
@@
We proceed by performing another change of variables, $w^{\pm}=\tilde{u}\pm v$, to diagonalize the remaining equation:
@@
Once we arrive at (17), it is relatively straightforward to construct an exponential dichotomy using the spectral decomposition of the spherical Laplacian and the finite spectral decomposition of $V_{\infty}$ (diagonal thanks to our hypothesis (H3)).
@@
Suppose that (H1), (H2), and (H3) all hold. Then equation (17) has an exponential dichotomy,  $\tilde{\Phi}^{\mathrm{s}/\mathrm{u}}(r,t)$, on $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ with the same decay rate as the asymptotic equation (6):  $\eta=\displaystyle{\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
@@
If (H1), (H2), and (H3) all hold, then (2) has an exponential dichotomy in $X_{r}$ for $r>1$ with the same decay rate as the asymptotic equation (6): ${\eta=\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
@@
In this section we will fill in the details of the argument by proving Lemma 2.4, Lemma 2.5, and Theorem 2.1. We begin by discussing, in detail, the spectral decomposition of the spherical Laplacian and the resulting multiplication symbols for the operators $A(r)$ and $A(r)^{1/2}$.
@@
For each degree $k\geq 0$ with $k\in\mathbb{Z}$, there is an orthonormal set $\{Y_{k}^{j}\}_{j=1}^{\nu(k)}\subset L^{2}(S^{n-1},\mathbb{C})$ which spans a $\nu(k)$-dimensional eigenspace and satisfies
@@
These $\{Y_{k}^{j}\}$ are the spherical harmonics of $\Delta_{S^{n-1}}$, and for each $(u(r,\cdot),v(r,\cdot))\in L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$ they allow a spectral decomposition given by
$u(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}u_{kj}(r)Y_{k}^{j}(\varphi)\quad\text{where}\quad u_{kj}(r)=\int_{S^{n-1}}u(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d}$
$\text{and}\quad v(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}v_{kj}(r)Y_{k}^{j}(\varphi)\quad\text{where}\quad v_{kj}(r)=\int_{S^{n-1}}v(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d}.$
This decomposition also allows for the definition of alternative, equivalent Sobolev norms on $H^{\alpha}(S^{n-1},\mathbb{C}^{d})$, and similarly on $L^{2}(S^{n-1},\mathbb{C}^{d})$:
$\|u(r,\cdot)\|_{H^{\alpha}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}(1+\mu_{k})^{\alpha}\|u_{kj}(r)\|^{2}_{\mathbb{C}^{d}},\qquad\|u(r,\cdot)\|_{L^{2}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}$ (18)
According to this decomposition, Equation (9) is reduced in each degree $k$ (suppressing the $j$ subscript as each equation is the same) to
$\begin{pmatrix}{u_{k}}^{\prime}\\ {v_{k}}^{\prime}\end{pmatrix}=\begin{pmatrix}0&1\\ V_{\infty}+\mu_{k}/r^{2}&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}=\begin{pmatrix}0&1\\ A_{k}(r)&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}$ (19)
@@
Notice that this spectral decomposition reveals $A(r)$ as a multiplication operator with symbol $A_{k}(r)$. Similarly, $A(r)^{1/2}$ can be considered a multiplication operator with symbol $A_{k}(r)^{1/2}$. When $d>1$, we can further decompose the matrices $A_{k}(r)$ and $A_{k}(r)^{1/2}$ in terms of the Jordan normal form of $V_{\infty}$. Recalling that by hypothesis (H3), $V_{\infty}$ is diagonalizable with eigenvectors $\{\rho_{1},\cdots,\rho_{d}\}$, let $R$ be the matrix whose columns are each $\rho_{\ell}$. Then both $A_{k}(r)$ and $A_{k}(r)^{1/2}$ are also diagonalizable with the same eigenvectors as $V_{\infty}$:
$A_{k}(r)$ $=R\begin{pmatrix}\lambda_{1}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\!\lambda_{d}\end{pmatrix}R^{-1}+\mu_{k}/r^{2}=R\begin{pmatrix}\lambda_{1}+\mu_{k}/r^{2}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\lambda_{d}+\mu_{k}/r^{2}\end{pmatrix}R^{-1}\quad\text{and}$
$A_{k}(r)^{1/2}$ $=R\begin{pmatrix}\sqrt{\lambda_{1}+\mu_{k}/r^{2}}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\sqrt{\lambda_{d}+\mu_{k}/r^{2}}\end{pmatrix}R^{-1}=R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}$
where $\gamma_{k\ell}(r):=\sqrt{\lambda_{\ell}+\mu_{k}/r^{2}}$ are the eigenvalues of $A_{k}(r)^{1/2}$. In the case of nontrivial Jordan blocks, a similar decomposition of the symbols $A_{k}(r)$ and $A_{k}(r)^{1/2}$ would still reveal each $\gamma_{k\ell}(r)$ to be an eigenvalue of $A_{k}(r)^{1/2}$ with multiplicity inherited from $V_{\infty}$. In fact, the invariant generalized eigenspace of each eigenvalue would be preserved though the exact generalized eigenvectors would not.
In the analysis that follows, we will use the eigenvector basis $R=[\rho_{1},\ldots,\rho_{d}]$ for $\mathbb{C}^{d}$ to decompose $A(r)$, $A(r)^{1/2}$, and other related operators into scalar pieces. In particular, we will use this decomposition and the equivalence of finite norms to bound the $L^{2}$ norm of $A_{k}(r)^{1/2}$ and related operators in the proof of Lemma 2.4. This looks like the following:
@@
$\left\|A(r)^{1/2}u\right\|_{L^{2}}^{2}=\sum_{k,j}$ $\left\|A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}=\sum_{k,j}\left\|R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}$
$\leq\sum_{k,j}\left\|R\right\|^{2}\left(\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\right)\left\|R^{-1}\right\|^{2}\left\|u_{kj}\right\|_{\mathbb{C}^{d}}^{2}\leq C\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\|u\|_{L^{2}}^{2}$ (20)
We are now ready to proceed with the proof of Lemma 2.4 keeping in mind both the bound given by Equation (20) and the technique used to derive it. Note that a similar decomposition, not into scalar pieces but into nilpotent blocks, should yield a parallel argument if we relax hypothesis (H3) to consider nontrivial Jordan blocks in the spectral structure of $V_{\infty}$.
3.2 Proofs of Lemmas 2.4 and 2.5.

### v2

@@ 1 Introduction
(Definition 5.2 in [22]) We say that a radial spatial evolutionary equation of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy in the $r$-dependent family of Banach spaces $X_{r}$ on a sub-interval $J\subseteq\mathbb{R}^{+}$ if there exist positive constants $C$ and $\eta$, two strongly continuous families of projections $P^{\mathrm{s},\mathrm{u}}(r)\in L(X_{r})$ with $P^{\mathrm{s}}(r)+P^{\mathrm{u}}(r)=I$, and two families of linear operators $\Phi^{\mathrm{s},\mathrm{u}}(r,\rho)\in L(X_{\rho},X_{r})$ such that the following statements are true:
    $\|\Phi^{\mathrm{s}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq C\exp({-\eta(r-\rho)})\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\geq\rho,\,\,r,\rho\in J.$
    $\|\Phi^{\mathrm{u}}(r,\rho)\mathbf{u}_{0}\|_{X_{r}}\leq C\exp({\eta(r-\rho)})\|\mathbf{u}_{0}\|_{X_{\rho}},\qquad r\leq\rho,\,\,r,\rho\in J.$
@@
    $\,\,P^{\mathrm{s}}(r)\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)=\Phi^{\mathrm{s}}(r,\rho)P^{\mathrm{s}}(\rho)$
@@
Throughout this paper, we also consider exponential dichotomies posed on the fixed space $L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$; we omit the standard definition in the context of a fixed Banach space, but the interested reader can find one in [17], noting that it is analogous to the definition above.
For completeness, we now also state the robustness of exponential dichotomies in the context of the $r$-dependent family of function spaces $X_{r}$.
@@
(Proposition 5.9 in [22]) Exponential dichotomies are robust under sufficiently small perturbations. More precisely, if a system of the form $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$ has an exponential dichotomy on some interval $J\subseteq\mathbb{R}^{+}$ in the $r$-dependent family of $X_{r}$ spaces with constants $C,\eta>0$, then for all $\varepsilon>0$ and $\tilde{\eta}\in(0,\eta)$ there are constants $\tilde{C},\delta>0$ such that the perturbed system
@@
$\|\mathcal{B}(r)\|_{L(X_{r})}\leq\delta,\qquad r\in J$
@@
has an exponential dichotomy on $J$ in the same family of spaces, $X_{r}$, with constants $\tilde{C}$ and $\tilde{\eta}$ and with projections $\varepsilon$-close to the projections of $\mathbf{u}_{r}=\mathcal{L}(r)\mathbf{u}$. Furthermore, if $\|\mathcal{B}(r)\|_{L(X_{r})}=O(1/r)$, then one can take $\tilde{\eta}=\eta$.
The comparable standard robustness result in the setting of a fixed Banach space is given as Theorem 1 in [17]. Throughout the paper, we will invoke robustness in both settings.
To apply Proposition 2.2, we must first establish the existence of an exponential dichotomy for the unperturbed or principal system, in this case (14). Making reasonable assumptions on $V_{\infty}$, the asymptotic system (11) will have an exponential dichotomy on $L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$. The main difficulty lies in applying robustness to extend the exponential dichotomy for (11) to one for (14) on the family of $X_{r}$ spaces. The majority of the proof is therefore dedicated to this first application of robustness. Once we have an exponential dichotomy for the principal system (14), we then apply robustness a second time to obtain an exponential dichotomy on the full system (6).
The second application of robustness—to prove an exponential dichotomy for (6) given an exponential dichotomy for (14)—is direct, assuming that $\tilde{B}(r)$ is sufficiently small, and in fact, decaying like $\mathcal{O}(1/r)$. This will hold if the potential function $V(r,\varphi)$ decays like $\mathcal{O}(1/r)$ to the asymptotic potential $V_{\infty}$. We therefore make the following hypothesis:
  1. (H1.)
    There exists a constant $C>0$ such that for $r\geq 1$, $\|V(r,\varphi)-V_{\infty}\|\leq C/r$.
@@
Turning to the existence of an asymptotic exponential dichotomy, we note that since $\mathcal{A}_{\infty}$ is autonomous and homogeneous across $S^{n-1}$, we can explicitly construct an exponential dichotomy for (11) via a spectral decomposition whenever the point spectrum of $\mathcal{A}_{\infty}$ does not intersect the imaginary axis $\mathrm{i}\mathbb{R}\subset\mathbb{C}$. This will hold whenever the eigenvalues of $V_{\infty}$ do not lie on the negative real line, $(-\infty,0]$, so we therefore make the following hypothesis:
  1. (H2.)
    For each eigenvalue $\lambda_{\ell}$ of $V_{\infty}$, $\lambda_{\ell}\in\mathbb{C}\backslash(-\infty,0]$.
@@
Finally, we also assume that $V_{\infty}$ has only semi-simple eigenvalues, and is therefore diagonalizable. This last hypothesis greatly simplifies the proofs to follow, allowing for an explicit spectral decomposition of $\tilde{\mathcal{A}}(r)$ into scalar pieces. In §3.1 we will carry out this decomposition explicitly and comment on the extendability of related arguments to the case of nontrivial Jordan blocks.
  1. (H3.)
@@
    $V_{\infty}$ is diagonalizable with eigenvalues $\{\lambda_{\ell}\}_{\ell=1}^{d}$ and corresponding eigenvectors $\{\rho_{\ell}\}_{\ell=1}^{d}\subset\mathbb{C}^{d}$.
@@
Hypotheses (H2) and (H3) allow us to fully characterize an exponential dichotomy for the asymptotic system (11). We must then understand the relationship between the principal operator $\tilde{\mathcal{A}}(r)$, in (12), and the asymptotic operator $\mathcal{A}_{\infty}$. To do this, we establish an isomorphism, or rather a family of isomorphisms, that can translate between the $X_{r}$ setting and the $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ setting. The basis of this family of isomorphisms, established in Lemma 2.3, is the differential operator defined by
@@
    The family of operators $A(r)^{1/2}:H^{1}(S^{n-1})\rightarrow L^{2}(S^{n-1})$, where $A(r)$ is given in (15), is well-defined over all values of $r\geq 1$. In addition, the family is uniformly bounded, strongly continuous, and continuously differentiable in $r$ for $r>1$, with
@@
    $C_{1}\left(\|A(r)^{1/2}u\|_{L^{2}}^{2}+\|v\|_{L^{2}}^{2}\right)^{1/2}\leq\|\mathbf{u}\|_{X_{r}}\leq C_{2}\left(\|A(r)^{1/2}u\|^{2}_{L^{2}}+\|v\|_{L^{2}}^{2}\right)^{1/2}$
@@
Lemma 2.3 establishes $(u,v)\mapsto(A(r)^{1/2}u,v)$ as an isomorphism from $X_{r}$ to $L^{2}\times L^{2}$. Applying the corresponding change of variables $\tilde{u}=A(r)^{1/2}u$ to the principal equation (14) we get the following equation posed on $L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$:
@@
We proceed by performing another change of variables, $w^{\pm}=\tilde{u}\pm v$, to diagonalize the leading order part of (16):
@@
Once we arrive at (17), it is relatively straightforward to explicitly construct an exponential dichotomy using the spectral decomposition of the spherical Laplacian and the diagonalization of $V_{\infty}$.
@@
Under hypotheses (H1), (H2), and (H3), (17) has an exponential dichotomy,  $\tilde{\Phi}^{\mathrm{s}/\mathrm{u}}(r,t)$, in $L^{2}(S^{n-1})\times L^{2}(S^{n-1})$ on the interval $r\geq 1$ with the same decay rate as the asymptotic system, (11):  $\eta=\displaystyle{\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
@@
Under hypotheses (H1), (H2), and (H3), (6) has an exponential dichotomy in $X_{r}$ for $r\geq 1$ with the same decay rate as the asymptotic equation (11): ${\eta=\tilde{\eta}=\min_{1\leq\ell\leq d}\mathrm{Re}[\lambda_{\ell}^{1/2}]}$.
Lemma 2.4 will be proven in §3.2 by applying robustness of exponential dichotomies in $L^{2}\times L^{2}$ (Theorem 1 in [17]). Then, Theorem 2.1 will be proven in §3.3 by using Lemmas 2.3 and 2.4 to apply robustness of exponential dichotomies in $X_{r}$ (Proposition 2.2).
@@
Before proving Lemma 2.3, Lemma 2.4, and Theorem 2.1, we begin by discussing, in detail, the spectral decomposition of the spherical Laplacian and the resulting multiplication symbols for the operators $A(r)$ and $A(r)^{1/2}$.
@@
For each degree $k\geq 0$ with $k\in\mathbb{Z}$, there is an orthonormal set $\{Y_{k}^{j}\}_{j=1}^{\nu(k)}\subset L^{2}(S^{n-1},\mathbb{C})$ which spans the $\nu(k)$-dimensional eigenspace of $\Delta_{S^{n-1}}$ and satisfies
@@
These $\{Y_{k}^{j}\}$ are the spherical harmonics of $\Delta_{S^{n-1}}$. For each $(u(r),v(r))\in L^{2}(S^{n-1},\mathbb{C}^{d})\times L^{2}(S^{n-1},\mathbb{C}^{d})$ they allow a spectral decomposition given by
$u(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}u_{kj}(r)Y_{k}^{j}(\varphi),\quad u_{kj}(r)=\int_{S^{n-1}}u(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d},$
$\quad v(r,\varphi)$ $=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}v_{kj}(r)Y_{k}^{j}(\varphi),\quad v_{kj}(r)=\int_{S^{n-1}}v(r,\varphi)\overline{Y_{k}^{j}(\varphi)}d\varphi\in\mathbb{C}^{d}.$
Note that, unlike the Laplacian posed on $\mathbb{R}^{n}$, the spherical Laplacian posed on $S^{n-1}$ can be treated analogously in the case of $n=2$ and $n\geq 3$. If $n=2$, $\mu_{k}=k^{2}$ with $Y_{k}^{1}(\theta)=\exp({+\mathrm{i}k\theta})$ and $Y_{k}^{2}(\theta)=\exp({-\mathrm{i}k\theta})$ for $k>0$. See [11] and the references therein.
This decomposition also allows for the definition of alternative, equivalent Sobolev norms on $H^{\alpha}(S^{n-1},\mathbb{C}^{d})$, $\alpha>0$, and similarly on $L^{2}(S^{n-1},\mathbb{C}^{d})$:
$\|u(r,\cdot)\|_{H^{\alpha}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}(1+\mu_{k})^{\alpha}\|u_{kj}(r)\|^{2}_{\mathbb{C}^{d}},\qquad\|u(r,\cdot)\|_{L^{2}}^{2}:=\sum_{k=0}^{\infty}\sum_{j=1}^{\nu(k)}\|u_{kj}(r)\|_{\mathbb{C}^{d}}^{2}.$
According to this Fourier-like decomposition, (14) can likewise be decomposed into $u_{kj}$ components (suppressing the $j$ subscript, as each equation is independent of $j$)
$\begin{pmatrix}\partial_{r}u_{k}\\ \partial_{r}v_{k}\end{pmatrix}=\begin{pmatrix}0&1\\ V_{\infty}+\mu_{k}/r^{2}&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}=\begin{pmatrix}0&1\\ A_{k}(r)&0\end{pmatrix}\begin{pmatrix}u_{k}\\ v_{k}\end{pmatrix}$,
@@
Notice that this spectral decomposition reveals $A(r)$ as a multiplication operator with symbol $A_{k}(r)$. Similarly, choosing the principal branch of the square root, we consider the operator $A(r)^{1/2}$ as a multiplication operator with symbol $A_{k}(r)^{1/2}$.
@@
When $d>1$, we can further decompose the matrices $A_{k}(r)$ and $A_{k}(r)^{1/2}$ into scalar symbols via the diagonalization of $V_{\infty}$ given by hypothesis (H3). To do this, let $R$ be the matrix whose columns are given by the eigenvectors, $\{\rho_{1},\cdots,\rho_{d}\}$, of $V_{\infty}$. Then both $A_{k}(r)$ and $A_{k}(r)^{1/2}$ are also diagonalizable with the same eigenvectors as $V_{\infty}$:
$A_{k}(r)=R\begin{pmatrix}\lambda_{1}+\mu_{k}/r^{2}&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\lambda_{d}+\mu_{k}/r^{2}\end{pmatrix}R^{-1},\qquad A_{k}(r)^{1/2}=R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}$
where $\gamma_{k\ell}(r):=\sqrt{\lambda_{\ell}+\mu_{k}/r^{2}}$, defined to be the principal square root, are the eigenvalues of $A_{k}(r)^{1/2}$.
In the proof of Lemma 2.3, we repeatedly use this Fourier-like decomposition to obtain $r$-dependent upper-bounds for the $L^{2}(S^{n-1})$ norm of various operators related to $A(r)^{1/2}$. For example, for $A(r)^{1/2}$ we estimate
@@
$\begin{split}\Big\|A(r)^{1/2}u&\Big\|_{L^{2}(S^{n-1})}^{2}\!=\sum_{k,j}\left\|A_{k}(r)^{1/2}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}=\sum_{k,j}\left\|R\begin{pmatrix}\gamma_{k1}(r)&&\\[-10.00002pt] &\!\!\!\!\ddots&\\[-10.00002pt] &&\!\!\!\!\gamma_{kd}(r)\end{pmatrix}R^{-1}u_{kj}\right\|^{2}_{\mathbb{C}^{d}}\\ \!&\leq\sum_{k,j}\left\|R\right\|^{2}\left(\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\right)\left\|R^{-1}\right\|^{2}\left\|u_{kj}\right\|_{\mathbb{C}^{d}}^{2}\leq C\sum_{k,j}\max_{1\leq\ell\leq d}|\gamma_{k\ell}(r)|^{2}\|u_{kj}\|_{\mathbb{C}^{d}}^{2}\end{split}$ (18)
where additional bounds on the scalar symbols $\gamma_{k\ell}(r)$ can then be applied.
Remark.
In the case of nontrivial Jordan blocks, a similar decomposition of the symbols $A_{k}(r)$ and $A_{k}(r)^{1/2}$ could be applied via the generalized eigenvectors of $V_{\infty}$. This decomposition would still reveal each $\gamma_{k\ell}(r)$ to be an eigenvalue of $A_{k}(r)^{1/2}$ corresponding to each distinct eigenvalue $\lambda_{\ell}$ of $V_{\infty}$. For any $\lambda_{\ell}$ associated with a nontrivial Jordan chain, $\gamma_{k\ell}(r)$ would inherit the generalized eigenspace, though it would not inherit the same generalized eigenvectors. Consequently, for each nontrivial Jordan block of $R^{-1}V_{\infty}R$, the matrix $R^{-1}A_{k}(r)^{1/2}R$ would have a corresponding block of the same dimension which could be written as the sum of an $r$-dependent nilpotent matrix and the identity matrix multiplied by $\gamma_{k\ell}(r).$
Extending our results to this setting would require slightly different approaches to the proofs of Lemma 2.3 and Lemma 2.4 in §3.2. For Lemma 2.3, employing a similar approach would require obtaining bounds on each nontrivial block analogous to the ones proven in the Appendix. Alternatively, the theory of primary matrix functions (see, Ch 6 of [12]) suggests a less direct approach to proving the first statement of Lemma 2.3, but more investigation is required. For Lemma 2.4, though we give an explicit construction that relies on the diagonalizable nature of $V_{\infty}$, a less explicit proof still relying on robustness in $L^{2}\times L^{2}$ would likely follow even in the case of nontrivial Jordan blocks.
3.2 Proofs of Lemmas 2.3 and 2.4.